# Status Report on the Chicago-Carnegie Hubble Program (CCHP): Three Independent Astrophysical Determinations of the Hubble Constant Using the James Webb Space Telescope

### v1

@@ 12 Comparisons With Previously Published Data
12.1 TRGB Comparison with Archival SHoES Data for NGC 4258
Archival JWST data from the SHoES Cycle 1 program (GO-1685 , PI: A. Riess) offers an important constraint on our TRGB calibration, taken with different filters than our own data. In the left panel of Figure 17, we show a versus () color-magnitude diagram for a halo region of the galaxy NGC 4258 based on archival JWST data from the SHoES team. The position of the TRGB is shown by the solid black horizontal line at F090W = 25.08 mag. The right panel shows the marginalized and smoothed luminosity function in black, and the Sobel edge-detection filter output is shown in blue. Detection of the tip in the left panel, and its quantitative measurement in the right panel, is unambiguous. The luminosity function was smoothed by 0.1 mag before its digital first derivative was output by the Sobel filter.
The JWST -band () data were derived from the sum of four 258-second NIRCam exposures; the four equivalent exposures were taken in parallel, directed to the beam-split long-wavelength channel of NIRCam. The Sobel response function has a measured width of 0.10 mag, within which there are 624 red giant branch stars contributing to the measurement. We conservatively adopt an uncertainty of 0.02 mag for the statistical uncertainty. A foreground -band galactic line-of-sight extinction of 0.02 mag, to the halo of NGC 4258, as derived by Anand et al. (2024a) is adopted here. This gives a reddening-corrected magnitude of 25.06 mag for the TRGB in NGC 4258. Adopting a geometric distance modulus of 29.397 0.032 mag (Reid et al., 2019) yields our zero-point calibration of MF090W = -4.336 0.02 (stat) 0.032 [sys] mag. This can be compared to the zero point of MF090W = -4.32 0.025 (stat) mag published by Newman et al. (2024), averaging over six nearby galaxies observed with JWST, using the same filter combinations as discussed here, but zeroed to previously published TRGB distances using an -band calibration of MI = -4.05 mag (Freedman, 2021). The SHoES team (Anand et al., 2024b) has also published an averaged zero point for the TRGB method giving MF090W = -4.362 0.033 (stat) 0.045 [sys] mag (with different applications of the method differing by up to 0.04 mag). In all cases, the zero-point agreement is very good: Anand et al. is -0.025 mag brighter than our calibration, while Newman et al. (2024) are +0.017 mag fainter. Taking the average of the three determinations would return our zero point to within 0.003 mag.
12.2 Comparison of Previously Published TRGB and JAGB Distances
In Figure 18, we compare measured TRGB distances with those from JAGB stars, including nearby galaxies measured with ground-based telescopes, and tabulated in Madore & Freedman (2024a, b); Lee et al. (2024b). There is strong consistency, with remarkably low scatter amongst these two distance scales, over a range of a factor of 50 in distance. These comparisons provide an additional external constraint on both the zero-point calibrations of the TRGB and JAGB methods, as well as their respective internal precisions.
@@ 12.3 Comparison of H0 Values: -band TRGB Distances Measured Using HST
In Freedman et al. (2019, 2020); Freedman (2021) we presented results from a HST program to measure -band TRGB distances based on observations of 20 SNe Ia located in a sample of nearby galaxies. In Freedman (2021) four geometric anchor distances were applied to the band TRGB distance scale: (1) Gaia parallaxes for Milky Way globular clusters (Cerny et al., 2020; Maíz Apellániz et al., 2021; Vasiliev & Baumgardt, 2021); (2) the detached eclipsing binary distance (DEB) to the LMC (Hoyt, 2023; Pietrzyński, 2019); (3) the DEB distance to the SMC (Hoyt, 2023; Graczyk et al., 2020) and (4) the maser distance to NGC 4258 (Reid et al., 2019). This calibration, based on four geometric anchors, resulted in a value of = 69.8 0.6 (stat) 1.6 [sys] km s-1 Mpc-1, a number that is in excellent statistical agreement with the JWST analysis presented in this current study.
Of the 20 SNe Ia in Freedman (2021) and the 10 galaxies in the current study with TRGB distances, there are six galaxies in common to both studies: M101, NGC 1365, NGC 4038, NGC 4424, NGC 4536 and NGC 5643. The differences in distance moduli (in the sense JWST -band minus HST -band) are +0.098, -0.028, -0.070, -0.073, -0.026, and +0.137, respectively, with a weighted average distance modulus offset amounting to only +0.017 mag 0.014 (error on the mean) mag.
@@
12.4 H0 With Time
While both the precision and the accuracy of the measurements of distances to galaxies have improved considerably in recent years, there are still reasons to remain open as to whether our currently estimated uncertainties truly reflect the total uncertainties. Shown in Figure 19 are all published values from the database maintained by Ian Steer (Steer (2020), Steer (2024, priv. comm.) (small dark blue dots) and the values of calibrated by Cepheids (larger red filled circles).
There are several points worth noting in Figure 19. First, the values of obtained from SNe Ia, as calibrated by Cepheids, have remained remarkably constant over a quarter of a century. Second, the Cepheid-based values of are systematically higher than the (mean, median or mode) of the distribution of other determinations (Steer et al. 2020, 2024). Third, there is no bimodal distribution of published values centered, respectively, around the oft-quoted values of 67 and 73 km/s/Mpc. Future studies will be required to unambiguously confirm the higher values that are based on Cepheid measurements alone or determine whether systematic errors ultimately prove to be the explanation.
@@ 13 Future Prospects
With JWST, we have the means of improving both the accuracy and the precision of the locally determined extragalactic distance scale, solidly providing a multiply calibrated and independently verified determination of current expansion rate of the universe, .
@@
Improving the Zero-Point (Anchor) Calibration: There are four galaxies at present for which geometric distances can be measured for calibrating astrophysical/stellar distance indicators suitable for determining the extragalactic distance scale. These include the LMC, NGC 4258, and the SMC, with quoted uncertainties of 1% (Pietrzyński, 2019), 1.5% (Reid et al., 2019), 2% (Graczyk et al., 2020), respectively, and an estimated range of 1-3%. The Milky Way parallax offset currently has a significant impact on derived parallaxes for Cepheids and the period-luminosity-metallicity relation (e.g., see Groenewegen, 2021; Owens et al., 2022, for a discussion of the current uncertainties). Future data releases for the Gaia satellite are forecast to provide a 1% parallax calibration for the Cepheid PL relation. In addition, accurate parallax measurements for Milky Way globular clusters will provide a zero-point calibration for the TRGB, again at a level of 1%.
@@
Increasing the Numbers of SN Ia Host Calibrators: On average, only about one new SN Ia, per year, is found within a distance accessible for follow-up discovery of Cepheids, TRGB or JAGB stars using HST. Thus the total sample of SN Ia host galaxies is not expected to increase significantly over the next decade. At present there are 37 galaxies for which HST Cepheid distances have been measured as part of the SHoES project. Ten supernova-host galaxies have been observed with JWST as part of this paper (GO-1995), and an additional 11 have been observed with JWST for programs GO-1685 and 2875. One galaxy is in common to both programs, NGC 5643, and both programs additionally have observations of NGC 4258. JWST GO-3055 includes observations in the outer regions of NGC 1404 and NGC 1380 with the filters and , optimal for TRGB and JAGB measurements. In the case of JWST, doubling the sample of galaxies observed will improve the precision of by . Beyond improving the statistical uncertainties alone, observations of more distant galaxies where crowding effects are more severe, will be important for constraining potential systematic uncertainties increasingly encountered as a function of distance.
@@
Photometrically Consistent and Wider Areal Coverage of Nearby Galaxies: The Nancy Grace Roman Space Telescope888https://roman.gsfc.nasa.gov/, due to be launched in 2027, will have 100 times the field of view (FOV) of HST, with comparable resolution and sensitivity, with wavelength coverage extending from 0.48 to 2.3 m. The large FOV will enable the mapping of entire disks and halos of nearby SNe Ia host galaxies, giving simultaneous measurements of the TRGB, Cepheids, and JAGB stars in single (massive) pointings. The simultaneous measurements will provide a consistent photometric calibration for all three methods.
@@
Future SN Surveys: Over the last few decades, an enormous amount of time and energy has gone into ground-based surveys for SNe Ia (e.g., Pantheon+). This database is the merger of observations from 18 different surveys (Scolnic et al., 2022), taken with different telescopes, instruments, calibrations, etc. The Vera C. Rubin Observatory999https://rubinobservatory.org/, which will begin taking data in 2025, is forecast to discover over 3 million supernovae in its first decade of operation. Follow-up spectroscopy to measure redshifts using other ground-based facilities will pave the way for a dataset that has a consistent photometric calibration, without the need to apply photometric offsets and/or corrections to heterogeneous samples, as is the case today.
@@
Additional Tests for Systematics: Importantly, future measurement of employing techniques completely independent of the local distance scale, accurate at the 1-2% level, will be essential for ruling out remaining systematic uncertainties in the local distance scale. An example is gravitational wave sirens (e.g., Chen et al., 2018), which, if more objects can be discovered, hold exciting promise.
@@
More Accurate Modeling to Standardize SNe Ia: A promising path to increasing precision and accuracy in , at once addressing Points 2 and 5, is improvements to models for standardizing SN Ia Hubble residuals (e.g. Boone et al., 2021; Stein et al., 2022). For these models, the SN Ia Hubble residuals can be standardized to 0.07-0.08 mag in distance using time-series spectrophotometry (compared to typical dispersions in the 0.15 mag range for light curve standardization approaches). The smaller intrinsic dispersion of such models can both significantly increase the amount by which statistical precision of tightens per SN Ia calibrator and, perhaps more importantly, also tighten the constraints on any as-yet unseen systematics in SN Ia standardization, which are more likely to disproportionately impact the measurement of due to the currently small number of calibrator SNe Ia.
@@ 14 Summary and Conclusions
The infrared sensitivity and high resolution of the JWST is providing a powerful new means of measuring the distances to nearby galaxies, and thereby enabling new and independent determinations of . In this paper, we have measured the distances to 10 nearby galaxies using three independent astrophysical distance indicators: Cepheids, the TRGB and JAGB/carbon stars. SNe Ia have previously been well observed in all of these galaxies. For the JAGB distance scale, the data analysis from the raw data frames through to determination of was carried out blind.
An inter-comparison of the galaxy-to-galaxy distances results in agreement with a combined scatter in each case (i.e., Cepheid versus TRGB, TRGB versus JAGB, JAGB versus Cepheid) of less than 4%. This agreement represents a remarkable improvement from recent decades. In the case of the JAGB and TRGB distances, the results are even more striking, with agreement at a level of less than 1%.
A significant result from the current study is that the two distance indicators (TRGB and JAGB) that are least affected by crowding/blending or reddening, and that are based on a single, consistent calibration from JWST/ NIRCam, are giving larger distances (and therefore result in a lower ) than for the Cepheids. This difference in measurement of the local distances is independent of the ultimate step taken in applying the distances to SNe Ia, and thereby determining in the unperturbed cosmic flow. These differences are pointing to systematics affecting one or more of the distances and need to be better understood. However, while they do not rule it out, the results presented here do not lend strong support to the suggestion that there is missing fundamental physics in the early universe. Only future data will settle this issue unambiguously. Currently, the distances obtained using the TRGB and JAGB methods agree extremely well, and these methods have the advantage that the inner disks of galaxies can be avoided in their application, unlike Cepheids, which are young and embedded in regions of dust and high source densities and high surface brightness.
@@
One of the largest remaining uncertainties in measuring the local value of using SNe Ia (e.g., SHoES or CCHP) is the small number of nearby, calibrating SN Ia host galaxies. This small sample is a consequence of the fact that SNe Ia are relatively rare, and as a result, there are few galaxies in the local volume containing SNe Ia that are also close enough to resolve Cepheids, TRGB or JAGB stars with HST. Moreover, these serendipitously available supernovae also appear to be amongst the intrinsically brightest. But, in being nearby and well observed, they also have the lowest uncertainties. More data will be required to ascertain if there is a systematic error in the distances to the more distant SN Ia galaxies. In addition, the sample needs to be enlarged to simultaneously address potential bias and decrease the statistical uncertainties. Increasing this sample will require more time; unfortunately only 1-2 of these SNe Ia occur each year. In the context of resolving the tension, it is critical to decrease both the systematic and statistical uncertainties.
@@
Finally, we summarize our results in Figure 20. It shows the good agreement between the two cosmological values from Planck (Planck Collaboration et al., 2020), DESI (DESI Collaboration et al., 2024), and the three local distance scale values (and their averaged value) from CCHP JWST. The accuracy of our measurements is dominated by the systematic uncertainty in the distance to the anchor galaxy, NGC 4258, which is currently estimated to be 1.5%. Based on our three methods, applied to the CSP sample of SNe Ia, our best estimate of is 69.96 1.05 1.12 km s-1 Mpc-1. Individually, the values for the three methods agree well, with only 4 km s-1 Mpc-1 separating the largest and smallest values. The value of from this study falls between that obtained from studies of the CMB, from BAO measurements calibrated by BBN, and the SHoES value, all without significant tension. Future work on the local distance scale will shed light on whether or not additional physics is needed beyond the CDM cosmological model.

### v2

@@ 14 Comparisons With Previously Published Data
14.1 Comparison of TRGB and -band Distances
In Figure 17, we compare measured TRGB distances with those from AGB/carbon stars measured in the -band, including nearby galaxies measured with ground-based telescopes and HST, and tabulated in Freedman et al. (2025). There is strong consistency, with remarkably low scatter among these two distance scales, over a range of a factor of 50 in distance. These comparisons provide an additional external constraint on both the zero-point calibrations of the TRGB and JAGB methods, as well as their respective internal precisions.
@@ 14.2 Comparison of H0 Values: -band TRGB Distances Measured Using HST
In Freedman et al. (2019, 2020); Freedman (2021) we presented results from a HST program to measure -band TRGB distances based on observations of 20 SNe Ia located in a sample of nearby galaxies. In Freedman (2021) four geometric anchor distances were applied to the band TRGB distance scale: (1) Gaia parallaxes for Milky Way globular clusters (Cerny et al., 2020; Maíz Apellániz et al., 2021; Vasiliev & Baumgardt, 2021); (2) the detached eclipsing binary distance (DEB) to the LMC (Hoyt, 2023; Pietrzyński, 2019); (3) the DEB distance to the SMC (Hoyt, 2023; Graczyk et al., 2020) and (4) the maser distance to NGC 4258 (Reid et al., 2019). Using these four geometric anchors, the calibration yielded = 69.8 0.6 (stat) 1.6 [sys] km s-1 Mpc-1, a result that is in excellent statistical agreement with the JWST analysis presented in this study.
Of the 20 SNe Ia in Freedman (2021) and the 10 galaxies in the current study with JWST TRGB distances, there are six galaxies in common to both studies: M101, NGC 1365, NGC 4038, NGC 4424, NGC 4536 and NGC 5643. The differences in distance moduli (in the sense JWST -band minus HST -band) are +0.073, 0.006, -0.036, -0.074, -0.041, and +0.167, respectively, with a weighted average distance modulus offset amounting to only +0.008 0.030 mag (error on the mean).
@@
14.3 The Run of H0 over the Past 25 Years
While both the precision and the accuracy of the measurements of distances to galaxies have improved considerably in recent years, there are still reasons to remain open as to whether our currently estimated uncertainties truly reflect the total uncertainties. Shown in Figure 18 are all published values from the database maintained by Ian Steer (Steer (2020), and updated in 2024, priv. comm.; small dark blue dots) and the values of calibrated by Cepheids (larger red filled circles).
There are several points worth noting in Figure 18. First, the values of obtained from SNe Ia, as calibrated by Cepheids, have remained remarkably constant for over a quarter of a century. Second, the Cepheid-based values of are systematically higher than the (mean, median or mode) of the distribution of other determinations (Steer et al. 2020, 2024). Third, there is no bimodal distribution of published values centered, respectively, around the oft-quoted values of 67 and 73 km/s/Mpc. Future studies will be required to unambiguously confirm the higher values that are based on Cepheid measurements alone, or determine whether systematic errors ultimately prove to be the explanation.
@@ 15 Future Prospects
With JWST, we have the means of improving both the accuracy and the precision of the locally determined extragalactic distance scale, solidly providing a multiply calibrated and independently verified determination of the current expansion rate of the universe, .
@@
Improving the Zero-Point (Anchor) Calibration: At present, there are four galaxies for which geometric distances can be measured for calibrating astrophysical/stellar distance indicators suitable for determining the extragalactic distance scale. These include the LMC, NGC 4258, and the SMC, with quoted uncertainties of 1% (Pietrzyński, 2019), 1.5% (Reid et al., 2019), and 2% in distance (Graczyk et al., 2020), respectively. The Milky Way parallax offset currently has a significant impact on derived parallaxes for Cepheids and the period-luminosity-metallicity relation (e.g., see Groenewegen, 2021; Owens et al., 2022, for a discussion of the current uncertainties). Future data releases for the Gaia satellite are forecast to provide a 1% parallax calibration for the Cepheid PL relation. In addition, accurate parallax measurements for Milky Way globular clusters will provide a zero-point calibration for the TRGB, again at a level of 1%.
@@
Increasing the Numbers of SN Ia Host Calibrators: On average, only about one new SN Ia, per year, is found within a distance accessible for follow-up discovery of Cepheids, TRGB or JAGB stars using HST. Thus the total sample of SN Ia host galaxies is not expected to increase significantly over the next decade. At present there are 37 galaxies for which HST Cepheid distances have been measured as part of the SHoES project. Ten supernova-host galaxies have been observed with JWST as part of this paper (GO-1995), and an additional 11 have been observed with JWST for programs GO-1685 and 2875. One galaxy is in common to both programs, NGC 5643, and both programs additionally have observations of NGC 4258. JWST GO-3055 includes observations in the outer regions of NGC 1404 and NGC 1380 with the filters and , optimal for TRGB and JAGB measurements. In the case of JWST, doubling the sample of galaxies observed will improve the precision of by a factor of . Beyond improving the statistical uncertainties alone, observations of more distant galaxies, where crowding effects are more severe, will be important for constraining potential systematic uncertainties increasingly encountered as a function of distance.
@@
Photometrically Consistent and Wider Areal Coverage of Nearby Galaxies: The Nancy Grace Roman Space Telescope101010https://roman.gsfc.nasa.gov/, due to be launched in 2027, will have 100 times the field of view (FOV) of HST, with comparable resolution and sensitivity, with wavelength coverage extending from 0.48 to 2. 3m. The large FOV will enable the mapping of entire disks and halos of nearby SNe Ia host galaxies, giving simultaneous measurements of the TRGB, Cepheids, and JAGB stars in single (massive) pointings. The simultaneous measurements will provide a consistent photometric calibration for all three methods. The European Space Agency’s Euclid111111https://www.esa.int/Science_Exploration/Space_Science/Euclid_overview satellite has a wide field of view of 0.67 deg2 and is already providing detailed, high-resolution CMDs of the nearest galaxies (Hunt et al., 2024) out to 5 Mpc. Although the resolution is insufficient to measure SN Ia host galaxies directly, comparison of the distances measured using TRGB and JAGB stars will provide valuable tests of measurement systematics.
@@
Future SN Surveys: Over the last few decades, an enormous amount of time and energy has gone into ground-based surveys for SNe Ia (e.g., Pantheon+). This database is the merger of observations from 18 different surveys (Scolnic et al., 2022), taken with different telescopes, instruments, calibrations, etc. The Vera C. Rubin Observatory121212https://rubinobservatory.org/, which will begin taking data in 2025, is forecast to discover over 3 million supernovae in its first decade of operation. Follow-up spectroscopy to measure redshifts using other ground-based facilities will pave the way for a dataset that has a consistent photometric calibration, without the need to apply photometric offsets and/or corrections to heterogeneous samples, as is the case today.
@@
Additional Tests for Systematics: Importantly, future measurement of employing techniques completely independent of the local distance scale, accurate at the 1-2% level, will be essential for ruling out remaining systematic uncertainties in the local distance scale. Examples include gravitational wave sirens (e.g., Chen et al., 2018) and gravitationally lensed supernovae (e.g., Grillo et al., 2024), which hold exciting promise if more objects can be discovered and their precision and accuracy improved.
@@
More Accurate Modeling to Standardize SNe Ia: A promising path to increasing precision and accuracy in , at once addressing Points 2 and 5, is improvements to models for standardizing SN Ia Hubble residuals (e.g. Boone et al., 2021; Stein et al., 2022). For these models, the SN Ia Hubble residuals can be standardized to 0.07-0.08 mag in distance using time-series spectrophotometry (compared to typical dispersions in the 0.15 mag range for light curve standardization approaches). The smaller intrinsic dispersion of such models can both significantly increase the amount by which the statistical precision of tightens per SN Ia calibrator and, perhaps more importantly, also tighten the constraints on any as-yet unseen systematics in SN Ia standardization, which are more likely to disproportionately impact the measurement of due to the currently small number of calibrator SNe Ia.
@@ 16 Summary and Conclusions
The infrared sensitivity and high resolution of JWST offer a powerful new means of measuring the distances to nearby galaxies, enabling independent determinations of . In this study, we measured new distances to 10 nearby galaxies using JWST observations of TRGB and JAGB/carbon stars, all of which have well-observed SNe Ia. The JAGB distance scale analysis, from raw data to determination, was conducted blindly. A comparison of the galaxy-to-galaxy TRGB and JAGB distances yielded an average difference of less than 1%. This level of consistency marks a significant advancement compared to recent decades.
@@
A primary uncertainty in local measurements using SNe Ia (e.g., SHoES or CCHP) stems from the limited number of nearby calibrating SN Ia host galaxies. This scarcity arises from the rarity of SNe Ia, and as a result, there are few galaxies in the local volume containing SNe Ia that are also close enough to resolve Cepheids, TRGB or JAGB stars with HST. Notably, many of these nearby supernovae are among the intrinsically brightest. To determine if a systematic error exists in distances to the more distant SN Ia galaxies, a larger sample is needed. Expanding this sample will mitigate potential bias and reduce statistical uncertainties. Unfortunately these nearby SNe Ia occur only every one to two years. Addressing the tension will require a decrease in both the systematic and statistical uncertainties.
@@
Our results, which are summarized in Figure 1, illustrate the agreement between our CCHP HST + JWST TRGB measurement applied to the CSP SN Ia sample and the cosmological values from Planck (Planck Collaboration et al., 2020) and DESI (DESI Collaboration et al., 2024). The accuracy of our measurements is dominated by the systematic uncertainty in the distance to the anchor galaxy, NGC 4258. Our current best estimate of is 70.39 1.22 (stat) 1.33 (sys) 0.70 [] km s-1 Mpc-1. This value falls between that obtained from studies of the CMB, from BAO measurements calibrated by BBN, and the SHoES value, all without significant tension. While our results show consistency with CDM, continued improvement to the local distance scale is essential for further reducing both systematic and statistical uncertainties.
Appendix A Three Additional Geometric Anchors for the JAGB Distance Scale
In the following we present a preliminary comparison of independent, geometry-based zero points for the JWST JAGB distance scale based on an interpolation between ground-based (2MASS) -band (1.24 m) data and space-based HST (F110W) data for the same host-galaxy JAGB luminosity functions. From above and from below, these two filters closely bracket the JWST/NIRCam F115W filter, allowing a direct transformation from the published ground-based JAGB calibrations onto the JWST system. The JAGB stars in both the Large and Small Magellanic Clouds are currently too bright for JWST to image them without saturating. Below we offer a first look at an indirect solution to this problem: by defining a photometric transformation from both ground-based and HST-based -band photometry to the JWST/NIRCam system.
Figure A1 is a near-infrared (NIR) color-magnitude diagrams for the total population of high-luminosity resolved stars in M33 from the time-domain study of Javadi et al. (2017). These are time-averaged data. Approximately 100 JAGB stars were identified and their marginalized luminosity function is shown in the right panel, where the mode of the JAGB population is found at J = 18.50 0.006 (error on the mean). In a second, independent study of M33 (Lee et al., 2022) the mode of the (single-epoch) JAGB luminosity function for the outer-most region (shown in the right-most panel of their Figure 5) is found to be at J(2MASS) = 18.50 mag, with a quoted error on the mean of ±0.01 mag. The agreement between these two, totally independent datasets is reassuring.
A third NIR dataset for M33 comes from the HST study of the inner disk of this galaxy undertaken by Williams et al. (2021) using WFC3-UV and -IR. In Figure A2 we show the NIR CMD for that dataset, where the J-band imaging was taken through their F110W filter, which is 0.14 m to the blue of the 2MASS -band filter that is centered at 1.24 m. Our target wavelength, that of the F115W filter used by JWST/NIRCam is 1.15 m. It is conveniently straddled by the two available datasets.
Figure A2 shows the HST CMD and JAGB luminosity function for M33. The mode of the JAGB population (color-selected, as shown in red) is found at J(F110W) = 18.82 0.03 mag. (132 JAGB stars are found within the 0.64 (2-sigma) limits of the Gaussian fit, leading to an error on the mean of 0.027 mag.) The mode is fainter than the 2MASS value by +0.32 mag for the extended outer-disk population of JAGB stars in M33. A simple, linear interpolation between these two wavelengths gives a zero-point for a ground-based to JWST transformation tailored specifically to the AGB population, thereby avoiding the need to consider a color term. That transformation is J(F115W) = J(2MASS) + 0.205 mag.
Madore & Freedman (2020) give 2MASS absolute magnitude calibrations of the JAGB in the LMC & SMC, as being M(2MASS) = -6.22 and -6.18 mag, respectively. These transform to M(JWST) -6.015 and -5.975 mag in the JWST/NIRCam F115W system. Additionally, we recall that a Milky Way (2MASS) absolute magnitude JAGB calibration, based on individual parallaxes, can be found in Madore et al. (2022b). Their averaged Milky Way value is = -6.19 0.04 (stat), transforming to = -5.985 0.04 mag
The global average of the three new (MW, LMC & SMC) geometric anchors and including the NGC 4258 direct calibration of = -5.99 0.15 mag, gives = -5.99 0.015 mag (stat). This is -0.005 mag (marginally) brighter (0.25% further in distance) than the value used in (Lee et al., 2024a), based on NGC 4258 alone, and presented again in the main text of this paper.
Appendix B Differences with Analyses in R24
Here we discuss a few points in response to those raised in R24.
1.
Method for correlated errors and H0:
The analysis in the current paper differs substantially from an earlier preprint. We have postponed the measurement of a combined value based on three methods until larger samples of JAGB and Cepheid distances are available. The main results presented here are based on a significantly expanded TRGB galaxy sample. By combining our JWST TRGB results with earlier HST measurements (F19, F20, F21) we have increased the sample from 10 galaxies to 20 galaxies, hosting 24 SNe Ia. We have tabulated results for the smaller JWST TRGB and JAGB samples individually, but note that these values will improve as future JWST observations become available. To combine the JWST TRGB and JAGB results, we have averaged the distances per host galaxy, rather than averaging final values as done previously. This approach ensures that covariances such as systematic errors from zero-point calibration are accurately propagated into the final ho error budget.
We note that our Cepheid analysis is still ongoing. In the disk fields where Cepheids reside, comparisons of different software packages (DAOPHOT and DOLPHOT) are showing significant photometric differences that require further investigation. These systematic differences are not present in the less crowded fields used for the TRGB analysis.
2.
Method for measurement Uncertainties in NGC 4258 and Distance Comparisons:
The 1.5% geometric measurement uncertainty for NGC 4258 was applied to each individual TRGB, JAGB and Cepheid measurement, and does not average down in a combined analysis. The 0.07 mag error for the Cepheids in our earlier work was a deliberate conservative estimate, reflecting ongoing uncertainties in the Cepheid measurements. While the R24 Cepheid distances are in accord with our TRGB distances, the differences among Cepheid measurements (including our JWST data and R24) are larger than those between the TRGB and JAGB distances.
3.
Linearity of Cepheid Distances:
As more distant galaxies are observed with JWST, the linearity of the HST distance scale will continue to be tested. For example, there is a 3 correlation between MB and (shown in Figure B1 below). Shown are the absolute peak B magnitudes, MB, for the SHoES sample of 42 SNe Ia as a function of distance modulus, all based on the Cepheid distance scale and SN Ia magnitudes from R22. All data points (blue and orange) are from the R22 data set. The orange points represent the subset of SNe Ia observed as part of the current study. Our JWST sample selection was based on proximity, to minimize potential systematics from crowding/blending effects.
A 3 trend is evident in Figure B1, in the sense that the supernova absolute magnitudes become fainter with increasing distance. The slope, measured using a Python orthogonal-regression algorithm, scipy.odr, is 0.082 0.023, with a T-statistic (the slope divided by its standard error) of 3.6. This correlation is not an artifact of being present on both axes. For instance, a positive + deviation would shift a point to the right (larger distance modulus) and upward (M = m – and M is negative), while a negative - deviation would shift it left and downward. Such deviations would not result in the observed slope. Future observations will determine whether this trend reflects a genuine distance-dependent effect.
We reiterate that four of the five galaxies compared in R24a, which concluded that crowding effects are ruled out at an 8.2- confidence level, are at distances 20 Mpc. However, 60% of the SHoES sample are at distances 20 Mpc. Therefore, future JWST data are essential to complete this test.